### v1

# Image-Text-Image Knowledge Transferring for Lifelong Person Re-Identification with Hybrid Clothing States

Abstract.
With the continuous expansion of intelligent surveillance networks, lifelong person re-identification (LReID) has received widespread attention, pursuing the need of self-evolution across different domains. However, existing LReID studies accumulate knowledge with the assumption that people would not change their clothes. In this paper, we propose a more practical task, namely lifelong person re-identification with hybrid clothing states (LReID-Hybrid), which takes a series of cloth-changing and cloth-consistent domains into account during lifelong learning. To tackle the challenges of knowledge granularity mismatch and knowledge presentation mismatch that occurred in LReID-Hybrid, we take advantage of the consistency and generalization of the text space, and propose a novel framework, dubbed $Teata$, to effectively align, transfer and accumulate knowledge in an “image-text-image” closed loop. Concretely, to achieve effective knowledge transfer, we design a Structured Semantic Prompt (SSP) learning to decompose the text prompt into several structured pairs to distill knowledge from the image space with a unified granularity of text description. Then, we introduce a Knowledge Adaptation and Projection strategy (KAP), which tunes text knowledge via a slow-paced learner to adapt to different tasks without catastrophic forgetting. Extensive experiments demonstrate the superiority of our proposed $Teata$ for LReID-Hybrid as well as on conventional LReID benchmarks over advanced methods.
Person Re-Identification, Lifelong Learning, Hybrid Clothing States
ccs: Computing methodologies Object identification
Figure 1. Conventional lifelong ReID task (LReID) assumes pedestrians would always exhibit the same clothing in the lifelong evolution process of ReID models. To promote ReID applications, we propose a more practical task by taking the clothing changes into account. It is much more challenging since the mismatch of knowledge caused by the hybrid clothing states across domains can lead to learning confusion and exacerbate knowledge forgetting. “SC” and “CC” denote same-cloth and cloth-changing states, respectively.
1. Introduction
Lifelong Person Re-identification (LReID) is a sub-field of person ReID, which aims to retrieve the same individual that appeared in different camera views, while incrementally accumulating the ReID knowledge in different scenarios (Pu et al., 2021). This topic offers numerous meaningful application scenarios, capable of accommodating the ongoing expansions of surveillance networks and facilitating continual learning of ReID models across diverse domains. Therefore, it has increasingly captivated attention and research efforts. For example, Ge et al. (Ge et al., 2022) replay old samples saved in memory with specific-designed supervisions to balance stability and plasticity for lifelong learning. Yu et al. (Yu et al., 2023) further introduce knowledge refreshing and consolidation techniques to achieve positive forward and backward transfer. To avoid data privacy, Lu et al. (Lu et al., 2022) present dream memory to preserve knowledge, while learning incrementally with augmented geometric distillation.
Despite advanced progress in these efforts, existing LReID approaches unfortunately ignore another factor that is common and of concern in the ReID application scenario, i.e., clothing changes (Gu et al., 2022; Qian et al., 2020; Jin et al., 2022; Yang et al., 2019). For example, models could be easily deceived by people with clothing changes (CC) even after sufficient steps of continual learning on conventional data, since the acquired knowledge has the bias that the same identity keeps the same clothes (SC). To overcome this limitation, we propose a more practical LReID task, namely lifelong person re-identification with hybrid clothing states (LReID-Hybrid), which requires constantly updating models on a series of cloth-changing and conventional ReID datasets. As depicted in Fig. 1, the goal of LReID-Hybrid is to improve the model’s ability to generalize across various scenarios with different clothing conditions.
However, developing a lifelong ReID model on sequential datasets with clothing state differences is a non-trivial task and inevitably encounters several challenges. The first one is knowledge representation mismatch. Although the incremental datasets all belong to the task of person ReID, they naturally have distribution differences due to locations and times, which causes visual representation shifts between different learning steps. This mismatch is much worse by changes in the state of people’s clothing. Clothing features are viewed as a discriminative representation in conventional ReID datasets, but this representation is disruptive when the same person changes their clothes (i.e., cloth-changing scenarios). Such a representation mismatch further exacerbates knowledge catastrophic forgetting, leading to fluctuations in the performance of the model during lifelong learning. The second challenge is knowledge granularity mismatch. In contrast to conventional ReID, cloth-changing ReID typically requires models to perceive more fine-grained cues in person images (Wang et al., 2023b), since visual appearance information, which takes up a greater proportion, can be conversely misleading. The difference in representation granularity is data-driven and has no impact on learning in a single data domain. Nevertheless, it occurs when carrying lifelong learning with multiple datasets, which inevitably hinders effective knowledge transfer and accumulation.
In light of the great success of visual language models in effective representation (Radford et al., 2021), as well as the advanced applications in ReID (Li et al., 2023), we seek a more compact and robust representation of person images in the text space. Natural language can intuitively and easily describe various features of person images, including appearance, gender, behavior, etc, some of which are conducive to the re-identification of person identity. More importantly, natural language can use a unified sentence pattern or vocabulary when describing person images, which can effectively bypass the domain differences existing in image pixels. Based on the above inspiration, we present a novel framework, dubbed Teata, that takes full advantage of Text modality to align, transfer and accumulate knowledge. Our core idea is a two-stage process of alternating iterations, to first distill pixel-level information and category labels into text semantic embeddings, and then project image features onto the text space to guide the image encoder to acquire ReID knowledge. Similar to CLIP-ReID (Li et al., 2023), we utilize CLIP technology to align image and text features, thereby transferring the accumulated knowledge in the text space.
To complement the text modality, we further propose the Structured Semantic Prompt (SSP) learning and the Knowledge Adaptation and Projection (KAP) strategy to tackle the aforementioned challenges. Concretely, we use learnable text prompts as a container for image-to-text distillation in the first stage, and decompose the prompt into several structured pairs. Each pair contains a set of tokens shared across all tasks, designed to learn representation objects of person images, and a specific set of tokens for learning identity-related image representation content. With the explicit structured design, we can transfer knowledge between different tasks with the uniform granularity of text descriptions. To alleviate the knowledge representation mismatch, we additionally tune the distilled text embeddings of different identities in the text semantic space during the second-stage learning. We optimize it with a slow-paced learner, so as to avoid knowledge confusion or forgetting and adjust for differences in representation between different tasks. Benefiting from the generalization of the text representation and the proposed novel modules, our Teata is rehearsal-free and capable of preserving old knowledge while adapting to new tasks for lifelong ReID of hybrid clothing conditions.
Contributions. We summarize the key contributions as follows,
(1) We exploit a more practical lifelong ReID task, namely LReID with hybrid clothing states (LReID-Hybrid). It requires incrementally learning a ReID model under different clothing states. To this end, we present a novel framework Teata to realize knowledge alignment, transferring and accumulation in an “image-text-image” closed loop.
(2) To alleviate the limitation of knowledge granularity mismatch, we propose the structured semantic prompt learning to explicitly decompose text prompts into a series of structured tokens. It unifies the granularity of text representation, thus facilitating the knowledge transfer.
(3) We introduce a knowledge adaptation and projection strategy to optimize text representations of every identity via a slow-paced learner. It can not only adapt to the representation differences of different tasks when guiding learning of image encoders, but also avoid knowledge forgetting concisely and effectively.
(4) Extensive experiments demonstrate the superiority of our proposed Teata for LReID-Hybrid over various baselines. Additionally, it is capable of achieving state-of-the-art results with significant margins on conventional LReID benchmarks.
2. Related Work
2.1. Person Re-Identification
There has been remarkable progress in person re-identification leveraging pre-prepared stationary training data (Wang et al., 2018; Luo et al., 2019; Zheng et al., 2019; Wang et al., 2023a). Conventionally, researchers focus on the same-cloth scenario where people would always show a relatively stable clothing appearance. Based on this assumption, to reduce the influence of intra-class variations, Zheng et al. (Zheng et al., 2019) propose to leverage the generated data and Zhou et al. (Zhou et al., 2019) propose to perform omni-scale feature learning. Wang et al. (Wang et al., 2018) design a multi-branch deep network architecture integrating discriminative information with various granularities. Some methods turn to auxiliary data clues (Jin et al., 2020; He et al., 2021) for help.
Although researchers have made efforts to promote the discriminative ability of ReID models, conventional person ReID models would fail when pedestrians change their clothes, since they almost exclusively focus on clothes. Xu et al. (Xu et al., 2021) and Eom et al. (Eom et al., 2021) attempt to use generative models to augment samples by explicitly synthesizing person images with various clothes, so as to learn more robust features against clothing changes. Some studies draw support from auxiliary modalities to assist learning soft-biometrics features unrelated to clothes (Qian et al., 2020; Yang et al., 2019; Wang et al., 2022; Jin et al., 2022; Hong et al., 2021; Chen et al., 2021; Liu et al., 2023; Guo et al., 2023; Agbodike et al., 2023; Cui et al., 2023). For instance, SPT+ASE (Yang et al., 2019) utilizes reliable and discriminative curve patterns on the body contour sketch. CAMC (Wang et al., 2022) uses heatmaps of human postures to encode body shape semantic information. GI-ReID (Jin et al., 2022) learns cloth-agnostic representations by leveraging personal unique and cloth-independent gait information. Other researchers take advantage of clothing labels to eliminate the negative effects of clothing features. For instance, CAL (Gu et al., 2022) adopts adversarial learning to penalize the model’s predictive power to clothes. AIM (Yang et al., 2023) adopts a dual-branch model to simulate causal intervention and eliminate clothing bias. However, simply ignoring the clothes of pedestrians would hurt the performance in the same-cloth scenario, in which clothing can be one of the most discriminative characteristics.
2.2. Lifelong Person Re-Identification
Lifelong learning seeks to maintain stable performance on old tasks while adapting the models to gain new knowledge (Li and Hoiem, 2017; Wu et al., 2019). Due to the expansion of smart surveillance systems, person ReID models also have to continually accumulate knowledge information from old domains and generalize well on new domains (Pu et al., 2021). To meet the demand, AKA (Pu et al., 2021) adopts knowledge distillation baselines (Li and Hoiem, 2017) to preserve acquired knowledge and maintains a learnable knowledge graph to adaptively update previous knowledge. However, due to significant domain variations, it struggles to retain old knowledge without access to previous data. GwFReID (Wu and Gong, 2021) formulates a comprehensive learning objective for maintaining coherence during progressive learning. Recently, rehearsal-based methods have achieved state-of-the-art results by saving a few exemplars in each domain. For instance, PTKP (Ge et al., 2022) proposes a pseudo task knowledge preservation framework to alleviate the domain gap in the last BN layer. KRC (Yu et al., 2023) introduces a dynamic memory model for bi-directional knowledge transfers and a knowledge consolidation scheme. However, due to privacy issues, it is impractical to save person images. As is opposite to reality, all previous works assume each pedestrian would always show a cloth-consistent appearance in the lifelong evolution process. They overlook the influence of clothing state variation and tend to fail when clothing information is not useful to distinguish different people. Differently, we explore a more practical lifelong setting where both cloth-consistent domains and cloth-changing domains are considered. We show that the consistency, compactness, and generalization of text semantics can help integrate different knowledge across domains with various clothing states.
Figure 2. Framework of our proposed Teata. It starts by aligning image features with text embeddings to transfer visual information and category labels to the text space. Next, we leverage the robustness and consistency of text semantics to guide the discriminative feature learning of visual representations. To achieve effective knowledge transfer, we design a Structured Semantic Prompt (SSP) learning in the first stage, which decomposes the original text prompt into several structured pairs to distill knowledge from the image space with a unified granularity of text description. Then, we introduce a Knowledge Adaptation and Projection strategy (KAP) in the second stage, which tunes text knowledge via a slow-paced learner to adapt to different tasks without catastrophic forgetting.
3. Methodology
3.1. Formulation and Overview
Task definition. In the task of lifelong person ReID with hybrid clothing states, a stream of datasets $\mathcal{D}={\{\mathcal{D}^{(t)}\}}^{T}_{t=1}$ are sequentially collected in different domains. $\mathcal{D}^{(t)}$ denotes the $t$-th added ReID dataset, which could be either the Same-Cloth scenario (i.e., each pedestrian always wears the same clothes in different camera views), or the Cloth-Changing setting (i.e., images of the same person can be captured appearing different clothing). It is a more practical yet challenging task than previous LReID (Ge et al., 2022; Pu et al., 2021; Yu et al., 2023), since the model is required to perform well on seen or unseen domains with various clothing conditions, pursuing the ability to not only defend the catastrophic forgetting problem, but also balance the learned cloth-related and cloth-irrelevant identity knowledge. At the $t$-th step, $\mathcal{D}^{(t)}=\{\mathcal{D}^{(t)}_{train},\mathcal{D}^{(t)}_{test}\}$ contains the training set and the testing set, each of which includes a variety of pairs of images $x_{i}^{(t)}$ and identity labels $y_{i}^{(t)}$. An image encoder $\mathcal{G}_{t-1}$ from the $(t-1)$-th step is thus trained on the training set, to incrementally acquire new knowledge on the $t$-th dataset, that is, $\mathcal{G}_{t-1}\to\mathcal{G}_{t}$. Notably, samples of the previous datasets are not available at the current step due to data privacy and storage limitations. Eventually, the ReID model $\mathcal{G}_{T}$ should perform well on all seen domains, which is evaluated on the testing sets $\{\mathcal{D}^{(t)}_{test}\}^{T}_{t=1}$ at the end of the lifelong learning process of totally $T$ steps. Meanwhile, $\mathcal{G}_{T}$ should generalize well on the testing sets of unseen domains $\{\mathcal{D}^{(u)}_{u\_test}\}^{U}_{u=1}$, where $U$ is the total number of unseen datasets with various clothing states.
Method overview. We propose a new framework Teata, to effectively align, transfer and accumulate knowledge using the text space as an intermediate medium. The overview of our framework is illustrated in Fig. 2. Inspired by CLIP-ReID (Li et al., 2023), we adopt CLIP (Radford et al., 2021) to bridge image and text latent space, and build a closed loop of “image-text-image”. Teata starts by aligning image features with text embeddings to transfer visual information and category labels to the text space. Next, we leverage the robustness and consistency of text semantics to guide the discriminative feature learning of visual representations. This text-to-image knowledge transfer is finally reflected in the parameter update of the image encoder.
Based on these two alternating stages, knowledge is continuously transferred in the image and text space, thereby achieving the accumulation of knowledge. To achieve effective knowledge transfer, we further design a Structured Semantic Prompt (SSP) learning in the first stage, which decomposes the original text prompt into several structured pairs to distill knowledge from the image space with a unified granularity of text description (see Sec. 3.2). To further alleviate the knowledge confusion caused by incremental learning from same-cloth ReID and cloth-changing ReID tasks, we introduce a Knowledge Adaptation and Projection strategy (KAP) in the second stage, which tunes text knowledge via a slow-paced learner to adapt to different tasks without catastrophic forgetting (see Sec. 3.3). Lastly, we describe the details of training and testing procedures of Teata in Sec. 3.4.
3.2. Structured Semantic Prompt Learning
CLIP-ReID (Li et al., 2023) utilizes a series of learnable tokens ${[X]}_{1}{[X]}_{2}\dots[X]_{M}$ as a container to distill visual knowledge from images. It is a free-from and task-driven learning process, as the features of each token learned could be different depending on different tasks or datasets. Although this token design is flexible, it could have a negative impact when performing lifelong learning across ReID datasets with different clothing states. Knowledge acquired from same-cloth datasets tends to focus on the visual appearance features, while it requires more fine-grained discriminative clues, such as shoe style, gender and body shape, when applying ReID models to the cloth-changing scenario. This mismatch, in terms of knowledge granularity, would cause a radically different update direction for each token, resulting in an inability to accumulate and transfer knowledge effectively.
Given a person image, we can describe its identity with text, for example, “a middle age man with short hair wearing a white skirt and black pants”. We observe that the description of attributes has a similar form of [content object]. Moreover, for instance, [short hair] is fine-grained information helpful for both same-cloth and cloth-changing ReID, and [white skirt] is appearance information that is more discriminative under the condition of changing clothes. It thus motivates us to present a Structured Semantic Prompt (SSP) with shared and specific tokens. Shared tokens are expected to function as [object] and be shared across all tasks so that the text description has similar granularity. The specific tokens represent [content] information unique to each identity, tailored for different ReID tasks. It is an ideal intuition, but unfortunately limited by the absence of attribute labels. To this end, we build upon learnable tokens and further structurally decompose them in the text semantics. More specifically, we assign the structured text prompt $P_{j}^{(t)}$, $j\in[1,N^{(t)}]$, to each pedestrian as follows.
(1) $P_{j}^{(t)}={\rm A\ photo\ of\ a}\ [{X}_{1}\ {Y}_{1}]\cdots[{X}_{M}\ {Y}_{M}]\ {\rm person.}$
where $M$ means the total number of token pairs, ${X}_{m}$ and ${Y}_{m}$ $(m\in[1,M])$ denote the specific and shared tokens, respectively. They are all learnable parameters with the same dimension and are alternately combined to form each structured pair $[{X}_{m}\ {Y}_{m}]$.
With our proposed structured semantic prompt, models are able to integrate knowledge across domains with various clothing states in the unified textual semantic space. Analyses in Sec. 4.4 and Tab. 6 further support its efficacy, albeit in a straightforward form. To optimize it, we adopt CLIP contrastive losses $\mathcal{L}_{i2t}$ and $\mathcal{L}_{t2i}$ (Radford et al., 2021) which constrain the bidirectional alignment of image features and text embeddings:
(2) $\mathcal{L}_{stage1}=\mathcal{L}_{i2t}+\mathcal{L}_{t2i}$
3.3. Knowledge Adaptation and Projection
Once the first stage of structured semantic prompt learning is done, we expect to transfer the knowledge from text to image space, so as to update the parameters of the image encoder $\mathcal{G}_{t}$. However, as shown in Fig. 2, it is noteworthy that we use the previous model of $\mathcal{G}_{t-1}$ to extract features of samples from the $t$-th dataset, and then distill knowledge to text embeddings. In the LReID-Hybrid task, the knowledge acquired from ReID tasks with different clothing states may be contradictory. For example, the visual appearance features are discriminate representations for same-cloth ReID, but may be misleading for cloth-changing ReID. This inevitably leads to learned text embeddings being sub-optimal or not applicable, thus affecting effective knowledge transfer and accumulation. To this end, we improve the second stage of learning with the Knowledge Adaptation and Projection (KAP) strategy.
Adaptation with slow-paced learner. After the first stage, text embeddings for each identity can be obtained using the learned structured semantic prompts and the frozen text encoder. Formally,
(3) $T^{(t)}={\left\{{f^{T}_{j}}^{(t)}=\mathcal{T}(P_{j}^{(t)})\ \big{|}\ j\in[1,N^{(t)}]\right\}}$
As discussed above, instead of freezing text embeddings $T^{(t)}$, we propose to fine-tune them with a slow-paced learner. It shows two benefits. First, we can continuously update text embeddings to make them more suitable for the current ReID task. Meanwhile, the slow-paced learner can prevent the learning of text embeddings from overfitting to the current task, resulting in forgetting previous knowledge. Note that when updating text embeddings, the text encoder $\mathcal{T}$ is kept frozen to reserve the learned text semantic knowledge and the structured semantic prompts are tuned for knowledge adaptation.
Project-then-learn. To transfer the knowledge from text to image space effectively, we use the text embeddings to guide the learning of $\mathcal{G}_{t}$ by projecting image features on the text space and then updating the ReID model, i.e., the image encoder, to accumulate new knowledge. Formally,
(4) $\mathcal{L}_{id}=-\sum\limits_{j=1}^{N^{(t)}}q_{j}\ \log\frac{\exp(f^{(t)}_{i}\cdot{{f^{S}_{j}}^{(t)}}^{\top})}{\sum\limits_{k=1}^{N^{(t)}}\exp(f^{(t)}_{i}\cdot{{f^{S}_{k}}^{(t)}}^{\top})}$
(5) $q_{j}=\left\{\begin{aligned} &1-\epsilon+\frac{\epsilon}{N^{(t)}}&&,j=y_{i}^{(t)}\\ &\frac{\epsilon}{N^{(t)}}&&,{\rm otherwise}\end{aligned}\right.$
where $f_{i}^{(t)}$ means the image feature of the $i$-th image with the identity label $y_{i}^{(t)}$, and ${f^{S}}^{(t)}$ is the adapted text embeddings and initialized with the L2 normalized text embeddings $T^{(t)}$, and ${f_{j}^{S}}^{(t)}$ denotes the adapted text embedding of the $j$-th identity. $N^{(t)}$ denotes the number of identities in the $t$-th dataset. $\epsilon$ is a small constant for label smoothing regularization (Szegedy et al., 2015), which is simply set to 0.1. Notably, Eq. 4 is effective in fine-tuning the image encoder $\mathcal{G}_{t}$ together with text embeddings. As a result, with the help of the slow-paced learner, both image encoder and text embeddings can be slowly updated so that new knowledge is accumulated without forgetting the old. We still keep the original projection loss $\mathcal{L}_{proj}$ proposed in (Li et al., 2023) to strengthen the learning of the image encoder, because it experimentally brings performance gains.
(6) $\mathcal{L}_{proj}=-\sum\limits_{j=1}^{N^{(t)}}q_{j}\ \log\frac{\exp(f^{(t)}_{i}\cdot{{f^{T}_{j}}^{(t)}}^{\top})}{\sum\limits_{k=1}^{N^{(t)}}\exp(f^{(t)}_{i}\cdot{{f^{T}_{k}}^{(t)}}^{\top})}$
where ${f_{j}^{T}}^{(t)}$ is the text embedding of the $j$-th identity.
Alternative variant. Benefiting from CLIP and our proposed SSP, text and image semantics are aligned in the shared latent space. Intuitively, image representations that are well aligned with text embeddings can also be used in Eq. 4 to guide knowledge transfer and accumulation. Specifically, we can average all the extracted image features of each identity to obtain image representations $V^{(t)}$ at the beginning of the second stage training. Formally,
(7) $V^{(t)}={\left\{{f^{V}_{j}}^{(t)}=\frac{1}{\left|X^{(t)}_{j}\right|}\sum\limits_{f^{(t)}\in X^{(t)}_{j}}f^{(t)}\ \bigg{|}\ j\in\left[1,N^{(t)}\right]\right\}}$
where $X^{(t)}_{j}$ denotes the set of all person images with the same identity label $j$, and |||*|| ∗ | denotes the cardinal number of the set. Then, L2 normalized $V^{(t)}$ is used for the initialization of the classifier in Eq. 4. In Sec. 4.4, we thoroughly investigate the efficacy of our designs, including knowledge adaptation with text embeddings (w/ KA-T), text-aligned image representations (w/ KA-V), random initialization (by default), and slow-paced learner (w/ SL).
3.4. Details of Training and Inference
Overall, two stages are involved in each step $t$ during the lifelong evolution of the model. In the first stage, both image encoder $\mathcal{G}_{t-1}$ and text encoder $\mathcal{T}$ and frozen, and Eq. 2 is adopted for the optimization of SSP. In the second stage, $\mathcal{G}_{t}$ is optimized with the guidance of the learned text embeddings. The losses involved in the second stage are formulated as follows.
(8) $\mathcal{L}_{stage2}=\lambda_{1}\mathcal{L}_{proj}+\lambda_{2}(\mathcal{L}^{b}_{id}+\mathcal{L}_{id})+\lambda_{3}(\mathcal{L}^{b}_{tri}+\mathcal{L}_{tri})$
wherein $\mathcal{L}^{b}_{id}$ and $\mathcal{L}^{b}_{tri}$ denote the identity classification loss (Luo et al., 2019) and triplet loss (Hermans et al., 2017) applied on the image features before the last projecting layer of the CLIP image encoder. $\mathcal{L}_{id}$ and $\mathcal{L}_{tri}$ are applied to the representations after projection. $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are coefficients to balance between different losses. After training in step $t$, only the image encoder $\mathcal{G}_{t}$ is used for ReID evaluation. The cosine distances between two person images as measured during inference.
4. Experiments
4.1. Datasets and Evaluation Protocol
Seen datasets. Two widely-used conventional same-cloth datasets, i.e., Market-1501 (Zheng et al., 2015) and MSMT17 (Wei et al., 2018), and two widely-used cloth-changing datasets, i.e., LTCC (Qian et al., 2020) and PRCC (Yang et al., 2019), are used in our proposed LReID-Hybrid benchmarks for training. The dataset statistics are shown in Tab. 1. We design six training orders to imitate different situations comprehensively as follows.
Order 1: Market-1501 $\rightarrow$ LTCC $\rightarrow$ MSMT17 $\rightarrow$ PRCC.
Order 2: Market-1501 $\rightarrow$ LTCC $\rightarrow$ PRCC $\rightarrow$ MSMT17.
Order 3: Market-1501 $\rightarrow$ MSMT17 $\rightarrow$ LTCC $\rightarrow$ PRCC.
Order 4: LTCC $\rightarrow$ Market-1501 $\rightarrow$ PRCC $\rightarrow$ MSMT17.
Order 5: LTCC $\rightarrow$ Market-1501 $\rightarrow$ MSMT17 $\rightarrow$ PRCC.
Order 6: LTCC $\rightarrow$ PRCC $\rightarrow$ Market-1501 $\rightarrow$ MSMT17.
Table 1. The statistical information of our LReID-Hybrid training benchmarks.
Dataset Names #IDs #images #cameras SC
Market-1501 (Zheng et al., 2015) 1,501 32,668 6 $\checkmark$
MSMT17 (Wei et al., 2018) 4,101 126,441 15 $\checkmark$
LTCC (Qian et al., 2020) 152 17,119 12
[…]
2613
PRCC (Yang et al., 2019) 221 33,698 3
[…]
2613
Table 2. Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 1). The results are reported after the last training phase. The best results are shown in bold.
Training Order 1 Market-1501 LTCC MSMT17 PRCC SC Average CC Average
mAP R-1 mAP R-1 mAP R-1 mAP R-1 $\overline{s}_{\rm mAP}^{s}$ $\overline{s}_{\rm R-1}^{s}$ $\overline{s}_{\rm mAP}^{c}$ $\overline{s}_{\rm R-1}^{c}$
AKA (Pu et al., 2021) 56.0 76.6 5.7 13.5 5.3 14.1 33.1 32.7 30.7 45.4 19.4 23.1
SFT 54.8 76.3 16.0 34.2 45.5 72.0 47.4 47.0 50.2 74.2 31.7 40.6
LwF (Li and Hoiem, 2017) 62.8 81.4 17.1 31.9 53.3 77.6 47.5 46.3 58.1 79.5 32.3 39.1
CLIP-ReID (Li et al., 2023) 61.0 81.2 16.8 33.7 44.5 72.2 47.3 46.1 52.8 76.7 32.1 39.9
Teata (Ours) 80.4 92.1 20.4 42.1 56.3 79.7 59.5 58.2 68.4 85.9 40.0 50.2
Joint-Train 89.1 95.0 13.6 28.8 70.3 86.6 45.8 47.6 79.7 90.8 29.7 38.2
We also evaluate our proposed method on the widely-used standard benchmarks involving only datasets without clothing changes (Wu and Gong, 2021; Ge et al., 2022), where Market-1501 (Zheng et al., 2015), DukeMTMC-reID (Zheng et al., 2017), CUHK-SYSU (Xiao et al., 2017) and MSMT17 (Wei et al., 2018) are used sequentially for lifelong learning. Note that DukeMTMC-reID is only used for academic use and fair comparison without identifying or showing pedestrian images. CUHK-SYSU is modified from the original for person search and rearranged following (Wu and Gong, 2021). For other datasets, we follow the training and evaluation protocols proposed in their original papers.
Unseen datasets. To evaluate the generalization of a model in the same-cloth scenario, we select five conventional ReID datasets, including CUHK01 (Li et al., 2013), CUHK02 (Li and Wang, 2013), GRID (Loy et al., 2010), SenseReID (Zhao et al., 2017) and PRID (Hirzer et al., 2011). We follow their standard protocols to conduct the query and gallery sets for testing. As for the cloth-changing scenario, we evaluate lifelong ReID models on VC-Clothes (Wan et al., 2020) and Celeb-reID (Huang et al., 2019), neither of which appeared in the training set. For the Celeb-reID dataset, we use its lightweight version (i.e., Celeb-reID-light) without losing generality, and each pedestrian in Celeb-reID-light does not wear the same clothes twice.
Evaluation metrics. We adopt mean Average Precision (mAP) and Rank-1 accuracy (R-1) for performance evaluation on seen domains so far after each training step. We also report the average accuracies by averaging mAP and R-1 in seen domains as well as the accuracies in unseen domains. Considering the great performance discrepancy between same-cloth and cloth-changing datasets and to evaluate the model’s ability in different clothing states, we report the average accuracies on seen same-cloth and cloth-changing datasets, respectively, which are denoted as $(\overline{s}_{\rm mAP}^{s},\overline{s}_{\rm R-1}^{s})$ and $(\overline{s}_{\rm mAP}^{c},\overline{s}_{\rm R-1}^{c})$. We denote the average accuracy on all unseen same-cloth and cloth-changing datasets as $(\overline{s}_{\rm mAP}^{us},\overline{s}_{\rm R-1}^{us})$ and $(\overline{s}_{\rm mAP}^{uc},\overline{s}_{\rm R-1}^{uc})$, respectively. For the LTCC, PRCC and VC-Clothes datasets, we evaluate in the cloth-changing setting where only cloth-changing samples are involved in the testing set.
4.2. Implementation Details
We adopt the CLIP (Radford et al., 2021) image encoder as the ReID model, which is ViT-B/16 (Dosovitskiy et al., 2021) with 12 Transformer (Vaswani et al., 2017) layers. The class token of the last layer is used as the visual representation of the input person image. The CLIP text encoder consists of 12 Transformer (Vaswani et al., 2017) layers and the [EOS] token of the last layer is considered as the text embedding given structured $P$ as input. The images are resized to 256 $\times$ 128 and augmented by random horizontal flipping, padding, cropping and erasing (Zhong et al., 2020). The batch size is set to 64, with 4 samples per pedestrian. For each domain, the model is optimized with Adam optimizer (Kingma and Ba, 2015) with weight decay of $1\times 10^{-4}$ in two stages. The first stage takes 120 epochs using a learning rate initialized as $3.5\times 10^{-4}$ with cosine learning rate decay. The second stage takes 60 epochs with the warmup strategy that linearly increases the learning rate from $5\times 10^{-7}$ to $5\times 10^{-6}$ in the first 10 epochs. The learning rate is then decreased by a factor of 10 at the 40th epoch. To achieve knowledge adaptation, we implement the slower-paced learner by decreasing the learning rate by a factor of 10 after the first domain. $M$ is set to 16. Following (Li et al., 2023), $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are set to 1, 0.25 and 1, respectively.
4.3. Comparison with the State-of-the-Art
Competitors. We first compare with a widely-used rehearsal-free method AKA (Pu et al., 2021) for LReID. Then, we reimplement all competitors using CLIP (Radford et al., 2021) pre-trained ViT-B/16 (Dosovitskiy et al., 2021) as the ReID model. Specifically, SFT denotes sequentially fine-tuning the model on each dataset, and Joint-Train denotes training on all datasets jointly. We also compare our proposed $Teata$ with LwF (Li and Hoiem, 2017) and CLIP-ReID (Li et al., 2023). LwF performs knowledge distillation loss on the logit values of old tasks to keep task-specific decision boundaries. CLIP-ReID leverages the CLIP (Radford et al., 2021) pre-trained text encoder and uses the learned text prompts to regularize the image encoder. CLIP-ReID is designed for conventional person ReID in a stationary domain, and we reimplement it to achieve lifelong learning by retraining text prompts on each sequentially coming dataset.
4.3.1. Results in Seen Domains
Without loss of generalization, we show the results of training with Order 1 and evaluating on seen datasets in Tab. 2. More results in other orders are provided in the appendix. We also show results on the standard LReID benchmarks in Tab. 3.
Table 3. Comparison with the state-of-the-art methods in the conventional LReID setting.w/ Ex.” denotes rehearsal-based methods using exemplars. The training order is Market-1501 $\rightarrow$ DukeMTMC-reID $\rightarrow$ CUHK-SYSU $\rightarrow$ MSMT17, and the results are reported after the last training phase. The best results are shown in bold.
Methods w/ Ex. Market-1501 DukeMTMC CUHK-SYSU MSMT17 SC Average
mAP R-1 mAP R-1 mAP R-1 mAP R-1 $\overline{s}_{\rm mAP}^{s}$ $\overline{s}_{\rm R-1}^{s}$
AKA (Pu et al., 2021) 59.7 80.1 32.7 48.3 82.0 84.4 17.1 34.9 47.9 61.9
GwFReID (Wu and Gong, 2021) $\checkmark$ 60.9 81.6 46.7 66.5 81.4 83.9 25.9 52.4 53.7 71.1
KRC (Yu et al., 2023) $\checkmark$ 60.3 82.3 58.7 72.7 88.9 90.5 43.3 67.7 62.8 78.3
PTKP (Ge et al., 2022) $\checkmark$ 75.8 89.7 62.0 76.7 85.0 86.3 34.5 60.9 64.3 78.4
LwF (Li and Hoiem, 2017) 54.0 76.2 64.7 80.7 85.6 87.5 57.7 80.1 65.5 81.1
CLIP-ReID (Li et al., 2023) 50.6 74.8 60.0 76.8 84.5 86.4 64.7 84.9 65.0 80.7
Teata (Ours) 81.0 91.8 70.7 82.1 94.0 94.8 67.4 85.7 78.2 88.6
Joint-Train 89.8 95.1 81.2 90.6 95.1 95.8 72.3 87.3 84.6 92.2
Figure 3. Comparison results of our proposed method with SFT on each seen domain along the lifelong learning process. We report the results in Order 1 w.r.t. (a) mAP and (b) R-1.
Results on LReID-Hybrid benchmarks. AKA (Pu et al., 2021) achieves relatively poor results in our proposed LReID-Hybrid task. Although AKA uses a knowledge graph to represent and accumulate knowledge, it still operates in the image space, making it hard to maintain and accumulate knowledge when different tasks exhibit significant differences, especially when clothing status changes. Due to the absence of data from previous domains, SFT suffers catastrophic forgetting greatly. LwF performs knowledge distillation between current and frozen old models to overcome catastrophic forgetting at the cost of sacrificing performance in the current domain. Due to the great discrepancy of acquired knowledge across domains caused by variations of clothing state, LwF cannot make a good balance between knowledge acquired from domains with different clothing states. Although leveraging text, CLIP-ReID (Li et al., 2023) has no obvious advantage compared with LwF which fine-tunes the pre-trained CLIP image encoder without the supervision of text. The results show that the potential of text is still not fully unleashed. On the contrary, our proposed method successfully takes advantage of text semantics to balance and accumulate knowledge from various domains and shows a significant superiority over them.
It is also worth noting the performance of Joint-Train is traditionally regarded as the proximate performance upper bound of the lifelong person ReID model. Similarly, when we train the standard and cloth-changing datasets jointly, we find the model can preserve its capability in the same-cloth scenario. However, the results in the cloth-changing domains are far from satisfactory. We argue that the acquired knowledge shows great discrepancy due to various clothing states across domains. The data imbalance in distribution, as well as the amount between standard and cloth-changing datasets, would make the model unexpectedly biased toward the former when they are used jointly for training. The results further reveal the challenges faced in our proposed LReID-Hybrid task that considers clothing variations during the real-world application of person ReID models. In contrast to the image model, the text encoder is trained by a mass of data, building a compact and well-represented space. Our proposed Teata leverages the text space as an intermediary for knowledge transfer and accumulation, which can effectively alleviate the representation bias caused by data imbalance.
In Fig. 3, we show the results in each seen domain along the lifelong learning process. Without the help of text semantics, SFT achieves inferior results in each domain. What’s worse, when adapting to new domains, it forgets previously acquired knowledge dramatically. Thanks to the consistent and compact text semantics, our proposed method can accumulate new knowledge while the loss of previously acquired knowledge is significantly reduced. We also observe a slight performance improvement on the LTCC dataset when training on PRCC at step 4 compared with step 3, and for our proposed method, the final R-1 at step 4 is even higher than that at step 2. We argue that since both PRCC and LTCC are cloth-changing datasets, the model tends to learn cloth-irrelevant identity representations on them. Therefore, the learned cloth-irrelevant identity clues on PRCC are useful in improving the discriminative ability of pedestrians with clothing changes on LTCC. Our proposed $Teata$ exhibits a smoother forgetting trend in performance, showing its effectiveness in transferring and accumulating knowledge.
Table 4. Unseen domain comparisons with state-of-the-art models trained in the LReID-Hybrid setting (Order 1). The results are given by the trained model after the last training step. $\dagger$’ means the light version of Celeb-reID. The best results are shown in bold.
Training Order 1 CUHK01 CUHK02 GRID SenseReID PRID VC-Clothes Celeb-reID SC Average CC Average
mAP R-1 mAP R-1 mAP R-1 mAP R-1 mAP R-1 mAP R-1 mAP R-1 $\overline{s}_{\rm mAP}^{us}$ $\overline{s}_{\rm R-1}^{us}$ $\overline{s}_{\rm mAP}^{uc}$ $\overline{s}_{\rm R-1}^{uc}$
SFT 71.9 70.7 63.9 65.5 23.0 15.2 53.8 46.2 46.3 34.1 33.9 33.9 8.3 17.2 51.8 46.3 21.1 25.6
Joint-Train 69.3 67.3 70.2 68.6 26.7 19.8 59.2 50.1 51.2 40.1 29.2 28.6 6.6 14.1 55.3 49.2 17.9 21.4
LwF (Li and Hoiem, 2017) 74.0 71.7 68.8 66.7 31.6 24.2 56.7 48.3 58.9 49.6 35.0 32.7 8.5 15.7 58.0 52.1 21.8 24.2
CLIP-ReID (Yu et al., 2023) 76.3 75.2 71.6 72.4 31.7 24.3 56.6 48.0 54.2 43.9 39.4 37.5 9.0 17.2 58.1 52.8 24.2 27.4
Teata (Ours) 84.0 82.6 77.0 75.3 49.6 41.6 62.5 54.2 73.8 65.0 41.0 40.2 17.7 35.3 69.4 63.7 29.4 37.8
Table 5. Unseen domain comparisons with state-of-the-art models trained in the conventional LReID setting. The training order is Market-1501 $\rightarrow$ DukeMTMC-reID $\rightarrow$ CUHK-SYSU $\rightarrow$ MSMT17. The results are given by the trained model after the last training step. $\dagger$’ indicates the light version of Celeb-reID. The best results are shown in bold.
Methods CUHK01 CUHK02 GRID SenseReID PRID VC-Clothes Celeb-reID SC Average CC Average
mAP R-1 mAP R-1 mAP R-1 mAP R-1 mAP R-1 mAP R-1 mAP R-1 $\overline{s}_{\rm mAP}^{us}$ $\overline{s}_{\rm R-1}^{us}$ $\overline{s}_{\rm mAP}^{uc}$ $\overline{s}_{\rm R-1}^{uc}$
Joint-Train 60.3 57.6 69.5 69.5 51.8 39.9 55.0 45.7 27.6 15.0 35.0 32.2 9.0 18.2 52.8 45.5 22.0 25.2
SFT 75.6 74.7 61.6 58.4 32.0 21.9 52.9 44.4 49.0 35.8 33.0 31.4 6.0 13.6 54.2 47.0 19.5 22.5
PTKP (Ge et al., 2022) 71.1 69.4 66.4 66.1 33.9 25.3 53.9 46.1 38.4 27.3 24.2 21.8 5.3 9.2 52.7 46.8 14.8 15.5
KRC (Yu et al., 2023) 79.1 78.0 70.0 69.0 33.6 25.5 51.3 44.0 51.8 40.6 20.2 17.5 6.3 12.2 57.2 51.4 13.3 14.9
LwF (Li and Hoiem, 2017) 76.5 74.7 67.2 66.7 36.8 28.0 56.2 47.0 58.0 46.4 35.0 33.5 6.6 13.2 58.9 52.6 20.8 23.4
CLIP-ReID (Yu et al., 2023) 77.2 75.6 65.3 64.9 35.4 25.3 56.7 48.5 53.3 40.5 36.4 34.9 6.0 11.4 57.6 51.0 21.2 23.2
Teata (Ours) 86.9 84.6 82.2 81.4 54.9 45.6 67.5 58.3 80.7 72.0 40.5 39.2 11.3 24.1 74.4 68.4 25.9 31.7
Results on cloth-consistent LReID benchmarks. Since our proposed method is capable of handling the challenging LReID-Hybrid task, it is expected to also work well when pedestrians do not change their clothes. As shown in Tab. 3, we also compare our proposed method with state-of-the-art lifelong person ReID methods on the standard LReID benchmarks. Exemplars are considered to be of vital importance in relieving the catastrophic forgetting problem, and replay-based methods (Wu and Gong, 2021; Yu et al., 2023; Ge et al., 2022) achieve better results than AKA (Pu et al., 2021) indeed. However, saving person images can cause privacy issues, which is not realistic in real-world applications, and can bring great resource consumption with the increase of training domains. We also implement LwF (Li and Hoiem, 2017) and CLIP-ReID (Li et al., 2023) on the cloth-consistent LReID benchmarks. With the superior capability of CLIP, they achieve better results than state-of-the-art replay-based methods (Wu and Gong, 2021; Yu et al., 2023; Ge et al., 2022) on the current domain, but the performance on previous domains is still far from satisfactory, showing the catastrophic forgetting problem is still not well dealt with. In contrast, our proposed method achieves a good balance across domains. With the help of consistent text semantics, our proposed method beats all competitors by a large margin without relying on impractical and inflexible exemplars.
4.3.2. Results in Unseen Domains
Furthermore, we discuss another important capability of the lifelong ReID model, its generalization ability in unseen domains, given the continuous accumulation of knowledge across different ReID tasks. In Tabs. 4 and 5, we report the results of models trained with Order 1 on the LReID-Hybrid task and trained on the standard LReID benchmarks in unseen domains, respectively.
Similar to the conclusions from the experiments in seen domains, Joint-Train performs better than SFT in the unseen same-cloth scenario but worse in the unseen cloth-changing scenario. The data imbalance in distribution, as well as the amount between datasets, make the model unexpectedly biased toward clothing-related knowledge when different clothing-state datasets are used jointly for training. Additionally, LwF (Li and Hoiem, 2017) and CLIP-ReID (Li et al., 2023) can improve the capability in unseen domains with the help of knowledge distillation and text modality. However, their performance is still far from satisfactory without dealing with knowledge representation mismatch and granularity mismatch properly. As expected, our method outperforms all other competitors on all unseen datasets, significantly showing its superior generalization ability. Thanks to our well-designed modules, $Teata$ can align, transfer and accumulate knowledge in an “image-text-image” closed loop, the consistent and generalized text semantics effectively promote the lifelong learning of ReID models, contributing to superior results in both seen and unseen domains.
Figure 4. Visualization of the learned representations in different domains with various clothing states. For $Teata$, we visualize the representations of (a) each image, and (b) each identity by averaging image representations with the same identity, respectively. We also visualize (c) the learned representations of each identity by SFT for comparison.
Secondly, we train models only on the conventional LReID task, but compare the results on both same-cloth and cloth-changing unseen domains, to explore in depth the generalization ability of lifelong ReID models. As shown in Tab. 5, we observe that (1) Our proposed $Teata$ beats all competitors by a large margin in different scenarios. Although it does not see any samples of changing clothes during the training phase, it is also capable of obtaining impressive results in the cloth-changing scenario. The results demonstrate the consistency and generalization of the text semantics, and that the designs of $Teata$ are effective in not only taking full advantage of text semantics, but also achieving knowledge accumulation in lifelong learning. (2) Since only the same-cloth datasets are used during training in the conventional LReID task, the advanced ReID methods (Ge et al., 2022; Yu et al., 2023) in Tab. 5 achieves inferior results in the unseen cloth-changing domains compare with those in Tabs. 4. (3) For the advanced ReID methods (Ge et al., 2022; Yu et al., 2023), with the performance improvement in the same-cloth scenario, the performance surprisingly decreases in the cloth-changing scenario. We explain that these models may exclusively focus on clothing information due to the overfitting of the training dataset bias, which is not reliable in the cloth-changing scenario. (4) LwF (Li and Hoiem, 2017) and CLIP-ReID (Li et al., 2023) are both built upon the pre-trained vision language model of CLIP. Benefiting from the information complementarity and generalization advantages of image-text multimodality, these two simple baselines show better results than previous lifelong ReID models (Ge et al., 2022; Yu et al., 2023). However, plagued by catastrophic forgetting due to knowledge representation and granularity mismatches, their performance is still sub-optimal. (5) Joint-Train shows no obvious advantage over other previous methods. Additionally, it is worth noting that it may work worse on some same-cloth datasets like CUHK01 (Li et al., 2013) and PRID (Hirzer et al., 2011), and unexpectedly works well in the unseen cloth-changing scenario. One possible reason is the mixture of data from different domains. Joint-Train tries to balance the domain gap, which inevitably sacrifices the performance of datasets with large domain deviations or small scale. On the other hand, clothing appearance in different domains will also appear the representation difference, which encourages the model to explore more robust representations rather than just clothing features. However, it still struggles with learning from different domains, resulting in limited performance.
4.4. Ablation Studies
Table 6. Ablation Studies of our proposed method. The average accuracies of four datasets in Order 1 at the last training step are reported.
No. SSP KAP SC Average CC Average
+KA-T +SL $\overline{s}_{\rm mAP}^{s}$ $\overline{s}_{\rm R-1}^{s}$ $\overline{s}_{\rm mAP}^{c}$ $\overline{s}_{\rm R-1}^{c}$
1 52.8 76.7 32.1 39.9
2 $\checkmark$ 58.3 80.7 33.8 42.8
3 $\checkmark$ $\checkmark$ 67.5 85.9 35.2 44.1
4 $\checkmark$ $\checkmark$ $\checkmark$ 68.4 85.9 40.0 50.2
Effectiveness of SSP and KAP. We propose SSP and KAP to tackle the knowledge mismatch challenges faced in the LReID-Hybrid task effectively. To handle the challenge of knowledge granularity mismatch, SSP transfers knowledge across the same-cloth and cloth-changing domains with the uniform granularity of text descriptions. As shown in Tab. 6, it effectively achieves performance improvement in both same-cloth and cloth-changing scenarios. Furthermore, after dealing with the knowledge representation mismatch problem via our KA-T and SL designs, our proposed $Teata$ achieves knowledge adaptation effectively, thus showing a significant performance improvement in Tab. 6.
Discussions of our KAP strategy. As discussed in Sec. 3.3, since image and text modalities are aligned in the shared latent space, it would also be effective to use the image representations to achieve knowledge adaptation. We verify it in Tab. 7 (“Ours w/ KA-V”) and observe similar results to ours. Furthermore, we try to remove the text alignment in the first stage and use the unaligned image representations for the classifier initialization (“Ours w/o Align.”). The great performance decrease confirms the effectiveness of the semantic alignment between image and text modalities. It suggests the advantage of leveraging the text space as an intermediary for knowledge transfer and accumulation across domains with hybrid clothing states.
Table 7. Ablation results of knowledge adaptation in our KAP strategy. The average accuracies of four datasets in Order 1 at the last training step are reported.
Methods SC Average CC Average
$\overline{s}_{\rm mAP}^{s}$ $\overline{s}_{\rm R-1}^{s}$ $\overline{s}_{\rm mAP}^{c}$ $\overline{s}_{\rm R-1}^{c}$
Ours w/o Align. 63.0 82.5 34.5 43.9
Ours w/ KA-V 69.0 86.1 39.5 49.2
Ours w/ KA-T 68.4 85.9 40.0 50.2
Visualizations of the learned representations. Our proposed $Teata$ aligns image and text features, thereby transferring the accumulated knowledge in the text space. Benefiting from the consistency, compactness and generalization of text semantics, the learned representations are compact across domains. To demonstrate it, we use t-SNE (Van der Maaten and Hinton, 2008) to visualize the learned representations in different domains. As shown in Fig. 4 (a), the learned image representations are distributed in a unified latent space regardless of domains. The identity representations also exhibit great compactness in Fig. 4 (b). Upon closer observation of Fig. 4 (b), the identity representations on the LTCC (Qian et al., 2020) and PRCC (Yang et al., 2019) datasets are distributed in a similar area. Since both of them involve clothing changes for each pedestrian, their identity representations show robustness to clothing variations, that is exhibit cloth-irrelevant semantics. A similar phenomenon happens on the Market-1501 (Zheng et al., 2015) and MSMT17 (Wei et al., 2018) datasets, where each identity shows consistent appearances in clothing. In contrast, as shown in Fig. 4 (c), the learned identity representations of SFT show great discrepancy for different domains. The learned identity representations on Market-1501 (Zheng et al., 2015), MSMT17 (Wei et al., 2018) and PRCC (Yang et al., 2019) are distributed separately. Due to the great variations in clothing on LTCC (Qian et al., 2020), the identity representations on which are dispersed. The inconsistent semantics can cause knowledge forgetting in the lifelong evolution process. Overall, with the effective text semantics, our proposed $Teata$ achieves the lifelong evolution of ReID models across domains and accumulates abundant knowledge to handle varying clothing states.
5. Conclusion
In this paper, we exploit a more practical task, namely lifelong person re-identification with hybrid clothing states (LReID-Hybrid), which takes a series of same-cloth and cloth-changing domains into account during the lifelong evolution of ReID models. We analyze the challenges of knowledge mismatch across domains with various clothing states and explore the effectiveness of text for LReID-Hybrid. We use CLIP to take advantage of the good alignment between image and text modalities. To complement the text modality, we further propose the Structured Semantic Prompt (SSP) learning and Knowledge Adaptation and Projection (KAP) strategy to tackle the challenges of knowledge mismatches in granularity and representation, respectively. Our proposed novel framework Teata realizes knowledge alignment, transferring and accumulation with an “image-text-image” closed loop. Extensive experimental results on both our proposed LReID-Hybrid and conventional LReID benchmarks demonstrate its effectiveness and superiority. Without any memory for replay, we show that leveraging consistent and generalized text semantics can already promote the performance of LReID significantly. We hope it can inspire more studies to focus on the lifelong evolution of ReID models with hybrid clothing states, promoting the landing of ReID systems.
Appendix
Appendix A More Comparison Results
To better verify the efficacy of our proposed framework Teata, we additionally show results in Orders 2$\sim$6 in Tabs. 8$\sim$12. We compare with one naïve solution of Sequential Fine-Tune (SFT), two rehearsal-free lifelong approaches AKA (Pu et al., 2021) and LwF (Li and Hoiem, 2017), as well as CLIP-ReID (Li et al., 2023) baseline that also uses text to promote ReID learning. Several similar observations can be made from tables that strongly demonstrate the effectiveness and robustness of our proposed method. (1) Teata is consistently superior to AKA, SFT and LwF in different training orders, outperforming them by a large margin. It clearly suggests the advantage of leveraging the text space as an intermediary for knowledge transfer and accumulation across different domains. (2) Although CLIP-ReID achieves relatively better results than other competitors in the same-cloth scenario, it degrades in the cloth-changing scenario. The reason for this phenomenon may be that i) the scale of same-cloth and cloth-changing datasets is unbalanced and ii) the knowledge representations for these two kinds of tasks are mismatched, which eventually leads to catastrophic forgetting, especially the knowledge for cloth-changing ReID. Thanks to our proposed designs of structured semantic prompt (SSP) learning and knowledge adaption and projection module, our Teata could successfully transfer and accumulate knowledge between same-cloth and cloth-changing domains, which bests CLIP-ReID with an overwhelming advantage. (3) As observed and discussed in Sec. 4.3 of our main paper, Joint-Train also performs better in the same-cloth scenario than in the cloth-changing scenario due to the discrepancy in different domains with various clothing states. On the contrary, Teata achieves a good balance in knowledge accumulation across domains, demonstrating the efficacy of our proposed method on lifelong learning.
Table 8. Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 2). The results are reported after the last training phase. The best results are shown in bold.
Training Order 2 Market-1501 LTCC PRCC MSMT17 SC Average CC Average
mAP R-1 mAP R-1 mAP R-1 mAP R-1 $\overline{s}_{\rm mAP}^{s}$ $\overline{s}_{\rm R-1}^{s}$ $\overline{s}_{\rm mAP}^{c}$ $\overline{s}_{\rm R-1}^{c}$
AKA (Pu et al., 2021) 52.2 75.0 7.6 17.6 26.5 25.8 21.8 41.8 37.0 58.4 17.1 21.7
SFT 49.5 74.5 11.6 25.3 22.3 20.8 63.8 84.7 56.7 79.6 22.8 23.1
LwF (Li and Hoiem, 2017) 58.7 79.6 14.2 27.6 32.1 32.0 60.2 82.1 59.5 80.9 23.2 29.8
CLIP-ReID (Li et al., 2023) 56.3 80.0 11.9 26.0 25.2 26.2 67.6 86.5 62.0 83.3 18.9 26.1
Teata (Ours) 82.9 93.0 17.7 35.2 47.4 45.9 65.9 85.0 74.4 89.0 32.6 40.6
Joint-Train 89.1 95.0 13.6 28.8 45.8 47.6 70.3 86.6 79.7 90.8 29.7 38.2
Table 9. Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 3). The results are reported after the last training phase. The best results are shown in bold.
Training Order 3 Market-1501 MSMT17 LTCC PRCC SC Average CC Average
mAP R-1 mAP R-1 mAP R-1 mAP R-1 $\overline{s}_{\rm mAP}^{s}$ $\overline{s}_{\rm R-1}^{s}$ $\overline{s}_{\rm mAP}^{c}$ $\overline{s}_{\rm R-1}^{c}$
AKA (Pu et al., 2021) 58.3 78.6 3.8 10.1 5.6 12.5 33.3 32.7 31.1 44.4 19.5 22.6
SFT 54.8 77.0 37.0 65.7 16.6 35.5 46.9 48.2 45.9 71.4 31.8 41.9
LwF (Li and Hoiem, 2017) 63.9 83.0 49.3 75.3 17.8 31.9 46.9 44.6 56.6 79.2 32.4 38.3
CLIP-ReID (Li et al., 2023) 63.6 82.7 45.9 73.7 18.8 40.1 49.1 48.7 54.8 78.2 34.0 44.4
Teata (Ours) 77.7 90.7 46.4 73.8 23.2 46.9 58.5 57.3 62.1 82.3 40.9 52.1
Joint-Train 89.1 95.0 70.3 86.6 13.6 28.8 45.8 47.6 79.7 90.8 29.7 38.2
Table 10. Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 4). The results are reported after the last training phase. The best results are shown in bold.
Training Order 4 LTCC Market-1501 PRCC MSMT17 SC Average CC Average
mAP R-1 mAP R-1 mAP R-1 mAP R-1 $\overline{s}_{\rm mAP}^{s}$ $\overline{s}_{\rm R-1}^{s}$ $\overline{s}_{\rm mAP}^{c}$ $\overline{s}_{\rm R-1}^{c}$
AKA (Pu et al., 2021) 11.4 25.0 27.9 51.3 29.6 32.6 11.5 27.1 19.7 39.2 20.5 28.8
SFT 11.5 25.8 53.7 78.1 24.2 23.9 63.5 84.4 58.6 81.3 17.9 24.9
LwF (Li and Hoiem, 2017) 13.6 25.0 71.3 87.9 31.7 29.1 55.5 79.4 63.4 83.7 22.7 27.1
CLIP-ReID (Li et al., 2023) 12.0 24.0 61.1 82.5 23.0 23.3 66.2 86.0 63.7 84.3 17.5 23.7
Teata (Ours) 17.4 37.2 72.4 88.9 48.6 49.6 64.9 83.9 68.7 86.4 33.0 43.4
Joint-Train 13.6 28.8 89.1 95.0 45.8 47.6 70.3 86.6 79.7 90.8 29.7 38.2
Table 11. Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 5). The results are reported after the last training phase. The best results are shown in bold.
Training Order 5 LTCC Market-1501 MSMT17 PRCC SC Average CC Average
mAP R-1 mAP R-1 mAP R-1 mAP R-1 $\overline{s}_{\rm mAP}^{s}$ $\overline{s}_{\rm R-1}^{s}$ $\overline{s}_{\rm mAP}^{c}$ $\overline{s}_{\rm R-1}^{c}$
AKA (Pu et al., 2021) 8.3 17.9 20.7 42.3 3.3 9.8 33.1 32.5 12.0 26.1 20.7 25.2
SFT 14.4 29.3 61.2 81.4 44.9 72.1 52.0 54.7 53.1 76.8 33.2 42.0
LwF (Li and Hoiem, 2017) 14.6 28.6 75.8 89.2 48.2 74.3 44.4 41.9 62.0 81.8 29.6 35.3
CLIP-ReID (Li et al., 2023) 14.3 33.2 62.3 81.7 41.7 70.0 42.5 42.3 52.0 75.9 28.4 37.8
Teata (Ours) 20.6 41.6 71.4 87.7 53.8 78.6 61.9 62.9 62.6 83.2 41.3 52.3
Joint-Train 13.6 28.8 89.1 95.0 70.3 86.6 45.8 47.6 79.7 90.8 29.7 38.2
Table 12. Comparison with the state-of-the-art methods in the LReID-Hybrid setting (Order 6). The results are reported after the last training phase. The best results are shown in bold.
Training Order 6 LTCC PRCC Market-1501 MSMT17 SC Average CC Average
mAP R-1 mAP R-1 mAP R-1 mAP R-1 $\overline{s}_{\rm mAP}^{s}$ $\overline{s}_{\rm R-1}^{s}$ $\overline{s}_{\rm mAP}^{c}$ $\overline{s}_{\rm R-1}^{c}$
AKA (Pu et al., 2021) 11.0 22.4 28.5 30.0 27.6 54.3 10.3 25.5 19.0 39.9 19.8 26.2
SFT 10.6 23.5 22.5 21.1 55.7 79.0 63.6 84.2 59.7 81.6 16.6 22.3
LwF (Li and Hoiem, 2017) 13.4 25.5 29.7 28.2 71.5 86.6 55.4 79.1 63.5 82.9 16.6 26.9
CLIP-ReID (Li et al., 2023) 11.7 23.5 26.2 25.4 61.9 83.2 65.7 85.3 63.8 84.3 19.0 24.5
Teata (Ours) 16.5 36.0 45.4 44.6 75.1 90.0 64.9 84.2 70.0 87.1 31.0 40.3
Joint-Train 13.6 28.8 45.8 47.6 89.1 95.0 70.3 86.6 79.7 90.8 29.7 38.2